OpenStudy (anonymous):

The sum of two consecutive numbers is 25. What is the largest of the consecutive numbers?

OpenStudy (anonymous):

15

OpenStudy (butterflydreamer):

The sum of two consecutive numbers is 25 So let the numbers be x, x+1 Because the question says the sum of two consecutive numbers That means x + (x+1) = 25 x+ x+ 1 = 25 2x +1 = 25 Subtract 1 from both sides! 2x = 25-1 2x = 24 Divide both sides by 2 x= 24/ 2 therefore x= 12 So the larger number would be (x+1) SOOOO 12+1 = 13 Your larger number would be 13 :)

OpenStudy (anonymous):

thank you

OpenStudy (butterflydreamer):

you're welcome